# Accelerated boundary integral analysis of energy eigenvalues for confined electron states in quantum semiconductor heterostructures

### v1

@@ 2.1. Boundary integral equations
where the exterior region is treated as an infinite domain, $P$ and $Q$ represent the source and field points, respectively, $c(P)$ stands for the solid angle coefficient [14], $\Gamma$ denotes the interface that separates the interior and exterior regions, the subscript $()_{,n}$ signifies the derivative taken with respect to the unit outward normal $\bm{n}=\bm{n}(Q)$ to $\Gamma$, e.g., $\psi_{,n}=\displaystyle{\partial\psi\over\partial\bm{n}}$, and $\mbox{d}Q$ refers to an infinitesimal boundary length.
@@ 2.2. Numerical implementation
If the interface boundary $\Gamma$ is discretized with $N$ boundary elements, Eqs. (3) and (4) transform to
$\sum_{k=1}^{N}G^{{\mathcal{I}}}_{k\ell}\psi^{{\mathcal{I}}}_{,n\ell}-\sum_{k=1}^{N}H^{{\mathcal{I}}}_{k\ell}\psi^{{\mathcal{I}}}_{\ell}$ $=$ 00 0 (11)
$\sum_{k=1}^{N}G^{{\mathcal{E}}}_{k\ell}\psi^{{\mathcal{E}}}_{,n\ell}-\sum_{k=1}^{N}H^{{\mathcal{E}}}_{k\ell}\psi^{{\mathcal{E}}}_{\ell}$ $=$ 00 0 (12)
By using the boundary conditions (10), Eq. (12) can be rewritten as
$\mu\sum_{k=1}^{N}G^{{\mathcal{E}}}_{k\ell}\psi^{{\mathcal{I}}}_{,n\ell}-\sum_{k=1}^{N}H^{{\mathcal{E}}}_{k\ell}\psi^{{\mathcal{I}}}_{\ell}=0$ (13)
If $\bm{u}_{{\mathcal{I}}}$ represents the vector encompassing all the unknown nodal degrees of freedom $\psi^{{\mathcal{I}}}$ and $\psi^{{\mathcal{I}}}_{,n}$ on the interface $\Gamma$, Eqs. (11) and (13) can be expressed as the following systems of linear equations:
@@
where
@@ 2.4. Accelerated energy eigenvalue search
By discretizing Eqs. (20) and (30) with $N$ boundary elements and using the first $(m+1)$ terms of the series expansion on the right hand side of these equations, the following system of equation is obtained:
$\sum_{j=0}^{m}\lambda_{{\mathcal{I}}}^{j}\Bigg{(}\sum_{k=1}^{N}\bar{G}^{{\mathcal{I}}}_{jk\ell}\psi^{{\mathcal{E}}}_{,n\ell}-\sum_{k=1}^{N}\bar{H}^{{\mathcal{I}}}_{jk\ell}\psi^{{\mathcal{E}}}_{\ell}\Bigg{)}$ $=$ 00 0 (36)
$\sum_{j=0}^{m}\lambda_{{\mathcal{E}}}^{j}\Bigg{(}\mu\sum_{k=1}^{N}\bar{G}^{{\mathcal{E}}1}_{jk\ell}\psi^{{\mathcal{E}}}_{,n\ell}-\sum_{k=1}^{N}\bar{H}^{{\mathcal{E}}1}_{jk\ell}\psi^{{\mathcal{E}}}_{\ell}+\eta\Big{(}\mu\sum_{k=1}^{N}\bar{G}^{{\mathcal{E}}2}_{jk\ell}\psi_{,n\ell}-\sum_{k=1}^{N}\bar{H}^{{\mathcal{E}}2}_{jk\ell}\psi_{\ell}\Big{)}\Bigg{)}$ $=$ 00 0 (37)
or in matrix form,
@@ 3.1. A quantum wire of circular cross-section
[…]
@@
[…]
@@ 3.2. A quantum wire of square cross-section
[…]
@@ 3.3. A quantum wire of rectangular cross-section
[…]
@@ 3.4. A quantum wire of oval cross-section
[…]

### v3

@@ 2.1. Boundary integral equations
[…]
@@ 2.2. Numerical implementation
If the interface boundary $\Gamma$ is discretized into $N$ boundary elements with a total of $N_{n}$ boundary nodes, collocating Eqs. (3) and (4) at each source point $P=1\ldots N_{n}$ respectively yields the following equations:
$\sum_{k=1}^{N_{n}}G^{{\mathcal{I}}}_{Pk}\psi^{{\mathcal{I}}}_{,nk}-\sum_{k=1}^{N_{n}}H^{{\mathcal{I}}}_{Pk}\psi^{{\mathcal{I}}}_{k}$ $=$ 00 0 (11)
$\sum_{k=1}^{N_{n}}G^{{\mathcal{E}}}_{Pk}\psi^{{\mathcal{E}}}_{,nk}-\sum_{k=1}^{N_{n}}H^{{\mathcal{E}}}_{Pk}\psi^{{\mathcal{E}}}_{k}$ $=$ 00 0 (12)
By using the boundary conditions (9) and (10), Eq. (12) can be rewritten as
$\mu\sum_{k=1}^{N_{n}}G^{{\mathcal{E}}}_{Pk}\psi^{{\mathcal{I}}}_{,nk}-\sum_{k=1}^{N_{n}}H^{{\mathcal{E}}}_{Pk}\psi^{{\mathcal{I}}}_{k}=0$ (13)
The collocation at all $N_{n}$ source points results in $N_{n}$ equations of the form (11) and another $N_{n}$ equations of the form (13). They are two systems of linear equations which can be written in matrix form as follows:
@@
where $\bm{A}$ and $\bm{B}$ are the coefficient matrices of order ($N_{n}\times 2N_{n}$),
@@ 2.4. Accelerated energy eigenvalue search
By discretizing Eqs. (20) and (30) with $N$ boundary elements having a total of $N_{n}$ nodes, and using the first $(m+1)$ terms of the series expansion on the right hand side of these equations, the following equations are respectively obtained for each source point $P=1\ldots N_{n}$:
$\sum_{j=0}^{m}\lambda_{{\mathcal{I}}}^{j}\Bigg{(}\sum_{k=1}^{N_{n}}\bar{G}^{{\mathcal{I}}}_{jPk}\psi^{{\mathcal{I}}}_{,nk}-\sum_{k=1}^{N_{n}}\bar{H}^{{\mathcal{I}}}_{jPk}\psi^{{\mathcal{I}}}_{k}\Bigg{)}$ $=$ 00 0 (36)
$\sum_{j=0}^{m}\lambda_{{\mathcal{E}}}^{j}\Bigg{(}\mu\sum_{k=1}^{N_{n}}\bar{G}^{{\mathcal{E}}1}_{jPk}\psi^{{\mathcal{I}}}_{,nk}-\sum_{k=1}^{N_{n}}\bar{H}^{{\mathcal{E}}1}_{jPk}\psi^{{\mathcal{I}}}_{k}+\eta\Big{(}\mu\sum_{k=1}^{N_{n}}\bar{G}^{{\mathcal{E}}2}_{jPk}\psi^{{\mathcal{I}}}_{,nk}-\sum_{k=1}^{N_{n}}\bar{H}^{{\mathcal{E}}2}_{jPk}\psi^{{\mathcal{I}}}_{k}\Big{)}\Bigg{)}$ $=$ 00 0 (37)
where Eqs. (9) and (10) are used in Eq. (37) to convert $\psi^{{\mathcal{E}}}_{,nk}$ and $\psi^{{\mathcal{E}}}_{k}$ into $\psi^{{\mathcal{I}}}_{,nk}$ and $\psi^{{\mathcal{I}}}_{k}$, respectively.
When applied to all $N_{n}$ source points, the two equations above form two linear systems, each containing $N_{n}$ equations. These systems can be represented in matrix form as follows:
@@ 3.1. A quantum wire of circular cross-section
Consider the problem of a GaAs quantum wire embedded in an infinite Ga0.63Al0.37As barrier region, which was previously studied in [12]. This wire features a circular cross-section with a radius of 5 nm. The parameters for this quantum structure are as follows: $m_{{\mathcal{I}}}^{*}=0.0665m_{0}$, $m_{{\mathcal{E}}}^{*}=0.0858m_{0}$, $V_{0}=320$ meV where $m_{0}$ denotes the free electron mass.
@@
Figure 2: Contour plots of the normalized probability density (in nm-2) for the two highest excited states in a quantum wire of circular cross-section (dashed line) at energy levels of (a) $E=188.551$ meV, and (b) $E=317.286$ meV.
@@ 3.2. A quantum wire of square cross-section
Figure 4: Contour plot of the normalized probability density (in nm-2) for the two highest excited states in a quantum wire with a square cross-section (dashed line) at energy levels of (a) $E=178.190$ meV, and (b) $E=274.008$ meV.
@@ 3.3. A quantum wire of rectangular cross-section
Figure 6: Contour plot of the normalized probability density (in nm-2) for the highest excited state in a quantum wire of rectangular cross-section (dashed line) at energy level of $E=273.176$ meV.
@@ 3.4. A quantum wire of oval cross-section
Figure 8: Contour plot of the normalized probability density (in nm-2) for the highest excited state in a stadium quantum wire (dashed line) at energy level of $E=184.373$ meV.